# SATA-Bench: Select All That Apply Benchmark for Multiple Choice Questions

### v2

@@ Abstract
Large language models (LLMs) are increasingly evaluated on single-answer multiple-choice tasks, yet many real-world problems require identifying all correct answers from a set of options. This capability remains underexplored. We introduce SATA-Bench, the first dedicated benchmark for evaluating LLMs on Select All That Apply (SATA) questions across diverse domains, including reading comprehension, law, and biomedicine. Our evaluation of 27 open-source and proprietary models reveals a significant gap: even the strongest model achieves only 41.8% exact match, exposing LLMs’ inability to reliably identify all correct answers. We find that this weakness stems from two core challenges: selection bias - models favor certain choices regardless of content, and count bias — models fail to predict the correct number of answers. To address these issues, we propose Choice Funnel, a decoding strategy that combines token debiasing with adaptive thresholding to guide models toward complete and accurate selections. Choice Funnel achieves up to 29% higher exact match than competitive baselines while reducing inference cost by over 64%. Our findings expose fundamental limitations in current LLMs and introduce a new framework for diagnosing and improving multi-answer reasoning. We release SATA-Bench and Choice Funnel to promote LLM development for robust decision-making in realistic, multi-answer applications.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable success across a variety of natural language processing tasks, with multiple-choice question (MCQ) answering emerging as a prominent evaluation setting [40, 56]. However, most LLM benchmarks and training pipelines focus on questions with a single correct answer among a fixed set of options (typically four answer choices). This design choice introduces a structural bias, limiting the models’ ability to generalize to more flexible, real-world tasks that require identifying multiple correct answers.
Many real-world scenarios demand such flexibility. For instance, a social media moderator might need to evaluate a post for several types of toxic content—such as threats, hate speech, or offensive language—and accurately tag all applicable categories. Similarly, journalists extracting information from news articles or biomedical researchers annotating scientific papers with multiple relevant subdomains face tasks that extend beyond single-choice frameworks. These are instances of Select All That Apply (SATA) questions, where more than four choices are presented and multiple answers are required. Existing LLMs often struggle in such scenarios by inaccurately selecting valid options. Figure 1 illustrates this bias and the shortcomings of even advanced LLMs.
Figure 1: This is a representative example to show that LLMs struggle with SATA (Select All That Apply) questions. The models often provide wrong answers when multiple correct options are present.
Despite the relevance of SATA-style tasks, there is a lack of standardized benchmarks and evaluation methods tailored to this setting. Existing MCQ benchmarks predominantly assess single-answer selection, leaving a gap in our understanding of how well LLMs perform in multi-answer scenarios. To address these challenges, we introduce SATA-Bench, a new benchmark suite specifically designed to evaluate and enhance LLM performance on SATA tasks. SATA-Bench is uniquely characterized by multiple correct answers to reflect real-world scenarios, and a diverse set of knowledge-based and reasoning-driven questions. By presenting LLMs with varying numbers of options and multiple valid answers, SATA-Bench enables the assessment of their ability and biases in more realistic scenarios. SATA-Bench extends the applicability of LLMs to diverse and complex real-world tasks.
@@
    SATA-Bench Data Curation: We curate a high-quality, diverse benchmark dataset explicitly designed to challenge LLMs on multi-answer tasks. SATA-Bench features 1,604 human-validated questions with varying difficulty levels, multiple correct answers, and carefully constructed distractor options. We provide readability, confusion, and similarity analyses to ensure clarity, diversity, and task complexity across six real-world domains: reading comprehension, news classification, event detection, toxicity identification, biomedical concept tagging, and legal document analysis.
@@
    Comprehensive Evaluation of 27 LLMs on SATA Tasks: We conduct the largest-to-date evaluation of 27 state-of-the-art proprietary and open-source LLMs on SATA questions, revealing that even top-performing models achieve a maximum exact match of only 41.8%. Our analysis highlights two key failure modes—selection bias and count bias—and demonstrates how current LLMs systematically underestimate the number of correct answers. We also break down performance across domains, showing that no single model dominates across all task types, underscoring the diversity and difficulty of SATA-BENCH.
@@
    Choice Funnel Decoding Algorithm: We propose Choice Funnel, a novel iterative decoding strategy that combines token debiasing and adaptive thresholding to mitigate selection and count bias. Choice Funnel outperforms three competitive baselines across 7 open-source models, achieving up to 29% improvements in exact match accuracy while reducing inference costs by over 64%. We demonstrate that Choice Funnel enables smaller models to rival or outperform much larger models on SATA tasks, offering a scalable solution for improving multi-answer reasoning.
@@ 2 SATA-Bench Data Curation
Our objective is to develop a dataset that encompasses a diverse range of tasks and domains and poses sufficient challenges to differentiate the capabilities of LLMs. The data curation process consists of three stages: selection of relevant source datasets, transformation of the data into SATA format, and filtering of questions for readability, diversity, and clarity (see Figure 2). We developed SATA-Bench questions to include tasks in Reading Comprehension [28], Text Classification (News [38], and Events [19]), and Domain Understanding (Toxicity [21], Biomedicine [41], and Laws [7]). Detailed descriptions of each source dataset are provided in Appendix A.
Table 1: Statistics of the SATA-Bench dataset (by data source). We report the following metrics: n: number of instances, LC: label cardinality, m: mean number of correct answers, me: median number of correct answers, min: minimum number of correct answers, max: maximum number of correct answers, r: ratio of the number of choices to the median number of correct answers (LC/me), w: mean word count, FRE: Flesch Reading Ease score, FGL: Flesch-Kincaid Grade Level score, ARI: Automated Readability Index, DCR: Dale-Chall Readability score, GFI: Gunning Fog Index, Confusion: mean confusion score. The final row summarizes these metrics across the entire SATA-Bench dataset.
Data Source n LC m me min max r w FRE FGL ARI DCR GFI Confusion
Reading Comprehension 328 3–15 2.8 2 2 10 na 2018.46 59.94 9.22 12.57 9.27 9.75 0.33
Toxicity 255 8 2.56 2 2 6 4 1015.32 37.83 12.28 13.33 10.49 12.57 0.27
News 248 6 2.36 2 2 5 3 785.93 62.51 8.92 11.15 11.1 10.94 0.26
Laws 311 15 5.3 5 2 10 3 5761.69 45.09 12.29 14.06 8.75 12.07 0.14
Events 202 6 2.63 2 2 5 3 3644.06 50.64 10.83 13.08 9.7 11.8 0.25
[…]
SATA-Bench 1604 3-16 3.55 3 2 10 3.2 2491.01 49.56 10.75 12.80 9.96 11.51 0.24
@@ 2.1 SATA Transformation
We transformed the original question/text to SATA questions following the steps below: 1. Collect text content, labels, and the total number of options for each question. 2. Maintain the option-to-answer ratio between 2 and 3 for consistency and improved difficulty [51], and determine the number of options ($k$) based on the number of correct answers for each question. 3. Generate distractor options for each question, consisting of the $c$ correct options and $k-c$ options randomly selected from the pool. 4. Shuffle the list of options to eliminate any label imbalance.
@@ 2.2 Question Filtering
Initial Filtering. To clean the original source data, questions with fewer than ten words were eliminated [44, 26]. Additionally, to ensure each question is understandable and solvable, we excluded those containing ambiguous, vague, or subjective terms (details are provided in Appendix B.1[33].
Readability. To ensure SATA-Bench questions are understandable and challenging, we assessed the readability of each question using the Flesch Reading Ease (FRE) score [20] and the Gunning Fog Index (GFI) [22]. We retained questions with an FRE score between 20-100 (inclusive), filtering out extremely easy or difficult questions [29], and a GFI score between 6-17, corresponding to 6th grade to graduate level difficulty [22]. This filtering removed unclear or trivial questions while maintaining a range of difficulties.  [^1] We also included four additional measures of readability (Flesch-Kincaid Grade Level (FGL) [29], Automated Readability Index (ARI) [29], and Dale Chall Readability (DCR) [12]) in the SATA-Bench dataset.
Question Similarity. Some of the original questions were too similar to each other, adding redundancy to the dataset. Following [58], we calculated the similarity score for each question pair using the cosine similarity score of the term frequency-inverse document frequency (TF-IDF) matrix [47] due to its efficiency. We removed questions sharing at least 80% cosine similarity to reduce duplication.
Confusion Score. The difficulty of SATA questions is largely related to how confusing the provided options are. We assessed the semantic similarity between the correct answers and the distractors as a measure of question-level confusion score [6]. We used ST5-XXL [36] for semantic similarity calculation, as it performed best in [34]. Then, to balance the confusion level, we grouped the questions into 10 bins according to the confusion score and sampled between 50-300 records from each bin to ensure SATA-Bench contains a diverse range of difficulty levels. Figures 5 and 7 show the distribution of the confusion scores before and after filtering as well as by source dataset. We release a dataset comprising 7,983 pre-validation questions as a by-product of our work.
Human Validation. Human evaluation was conducted in two stages. First, we used it to identify and remove questions containing ambiguous information as detailed in Appendix B.2. In the final stage, three human annotators reviewed all remaining questions to correct labeling errors; questions lacking consensus among all three annotators were excluded, as detailed in Appendix B.3. Statistics of the final SATA-Bench is shown in Table 1.
@@ 2.3 SATA-Bench Characteristics
After these curation steps, SATA-Bench has the following characteristics: (i) Multiple choices with multiple correct answers. All questions have multiple choices and more than one correct answers. (ii) Diversity. The questions come from various disciplines, including knowledge-based tasks such as domain understanding and reasoning-driven questions such as reading comprehension (see Figure 3). (iii) Human validation. All questions are manually validated to ensure they are clear and correct by readability scores and human validation. SATA-Bench removes ambiguous or overly simplistic items while maintaining a range of reading levels (from 6th grade to graduate level) (see Figure 3). (iv) Challenging. We constructed the SATA-Bench benchmark so that 76% have a FRE score within the standard range (60-70) and the average GFI score is approximately 13th grade (equivalent to the first year of college/university). The mean semantic similarity between correct answers and distractors (incorrect options) is 0.24, exhibiting a right-skewed distribution (skewness = 1.8). Most questions cluster around a similarity score of 0.22, with a few more difficult questions extending the tail towards the higher end (see Figure 3).
Figure 3: SATA-Bench Dataset Overview. SATA-Bench covers a diverse set of topics and achieves a balance between readability and difficulty (measured by confusion score). d1: Reading Comprehension, d2: Toxicity, d3: News, d4: Biomedicine, d5: Laws, and d6: Events.
@@ 3 Experiments
This section presents the experiments conducted to assess the capabilities of LLMs on SATA questions. The benchmark includes 16 proprietary models and 14 open-source models. (See Table 7 for full model cards.)
Experimental Setup. Because our benchmark contains diverse questions, we use a zero-shot evaluation. The system prompt specifies that each question has at least two correct answers, and we instructed the LLM to output the labeled result in JSONL format [25, 57]. Furthermore, we utilize a CoT prompting strategy as described in [37]. We then extract the answer from the JSONL file using both exact match and fuzzy match. For cases where JSONL extraction fails (fewer than 3% of cases), we use Claude 3 Haiku and Human Labelers to extract the correct options from the answers provided. For smaller models, the percentage of cases where JSONL extraction fails exceeds $5\%$, making the above methods less reliable. Following  [24], we remove CoT and use the probability of the first output token to retrieve options. We hold out a dataset of 100 randomly sampled instances from the benchmark dataset to generate a threshold for each model with optimal Jaccard Index [5]. We select all options with a probability greater than that threshold value. Note that this method applies only to models with accessible token probabilities. We have also included the performance of non-expert humans on the benchmark (see Appendix E).
Metrics. We evaluate models using metrics across three categories: performance, selection bias, and count bias (details in Appendix F). Performance is measured using Exact Match (EM), Jaccard Index (JI), Mean Average Precision (Precision), and Mean Average Recall (Recall). Selection bias includes RStd [56] and RSD [11, 42]. We also introduce Selection Probability Divergence (SPD) to quantify unselection bias, a form of selection bias where models consistently avoid certain options. Count bias is assessed using the mean count difference (CtDif), mean absolute count difference (CtDifAbs), and count accuracy (CtAcc).
Table 2: Performance comparison of 27 different LLMs across various metrics on SATA-Bench. We highlight the best (bold) and second-best (underline) values. Columns labeled [($\uparrow$)] indicate higher-is-better; columns labeled [($\downarrow$)] indicate lower-is-better. Models with explicit reasoning capabilities are highlighted in italic. All numeric values are rounded to two decimal places. We retrieve exact labels for models evaluated using Inference-Based Retrieval + CoT prompting. For models evaluated under Probability-Based Retrieval, we select labels based on token probability thresholds.
@@
[…]
O3 41.77 87.50 81.22 73.91 0.38 6.79 0.06 -0.39 0.94 46.12
GPT4.1 40.49 85.52 85.66 75.23 0.13 5.98 0.06 -0.04 0.85 45.52
Grok 3 Think 39.71 83.93 86.31 74.40 0.30 6.26 0.07 0.06 0.93 44.24
GPT4 39.47 85.90 83.17 74.11 0.21 6.63 0.06 -0.20 0.82 46.61
Claude 3.7 Think 37.92 85.03 78.77 70.96 0.46 18.77 0.34 -0.32 0.87 44.48
Claude 3.7 37.82 85.35 77.15 70.98 0.49 6.59 0.25 -0.43 0.93 43.58
Claude 3 Sonnet 36.49 84.58 78.81 70.72 0.36 7.37 0.07 -0.35 0.83 48.00
Geimini 2.5 Think 36.46 84.58 83.25 72.58 0.12 4.76 0.06 -0.01 0.88 43.76
Claude 3.5 Haiku 35.89 80.26 85.08 71.12 0.33 7.31 0.35 0.18 1.01 42.61
Claude 3 Haiku 35.64 83.59 80.16 70.63 0.42 6.24 0.07 -0.22 0.85 47.15
Claude 3 Opus 35.59 86.97 77.19 70.15 0.62 8.26 0.07 -0.52 0.93 44.36
Gemini 2 Flash 34.60 85.01 79.98 70.71 0.17 6.14 0.06 -0.23 0.91 39.94
GPT 4.1 mini 33.46 86.05 78.23 69.90 0.30 6.69 0.06 -0.39 0.97 38.61
Claude 3.5 Sonnet 32.22 87.57 75.25 67.15 0.43 8.41 0.09 -0.46 1.06 38.55
Llama 3.1 405B 30.17 86.24 75.31 67.18 0.33 6.90 0.45 -0.39 1.02 36.30
Deepseek R1 28.17 84.62 72.36 64.49 0.94 17.44 0.03 -0.57 1.13 33.52
Mistral Large V2 22.83 88.20 62.59 57.16 1.33 10.89 0.12 -1.10 1.47 27.27
Qwen Plus 21.12 88.54 59.53 55.74 2.24 10.72 0.11 -1.18 1.43 24.85
Llama 3.2 90B 18.30 89.56 60.80 55.78 1.84 11.10 0.27 -1.09 1.41 24.30
Llama 3.1 70B 17.94 89.56 60.64 55.59 1.81 10.06 0.10 -1.12 1.48 22.12
Non-expert Human 17.93 60.62 54.44 45.02 1.46 15.32 1.46 -0.6 1.44 34.12
@@
[…]
Mistral 8B 14.73 81.46 53.23 46.63 11.42 19.47 1.27 -1.35 1.95 21.01
Llama3 8B 13.82 80.30 47.37 43.64 12.09 17.85 1.09 -1.59 1.88 22.00
Bloomz 7B 11.27 66.09 50.80 41.15 20.62 29.00 1.51 -0.87 1.71 20.09
DeepSeek R1 Distill 8B 8.85 72.20 45.81 40.02 13.38 21.62 1.14 -1.29 1.75 20.42
Qwen2.5 14B 6.30 87.84 38.76 37.58 21.01 18.02 1.06 -2.24 2.26 11.93
Phi3 7B 2.97 87.25 35.67 34.57 23.22 18.57 1.22 -2.33 2.35 7.22
Phi4-mini-reasoning 2.12 77.98 30.82 29.69 21.62 13.90 1.59 -2.37 2.39 7.35

### v3

@@ Abstract
Current large language model (LLM) evaluations primarily focus on single-answer tasks, whereas many real-world applications require identifying multiple correct answers. This capability remains underexplored due to the lack of dedicated evaluation frameworks. We introduce SATA-Bench, a benchmark for evaluating LLMs on Select All That Apply (SATA) questions spanning six domains, including reading comprehension, legal reasoning, and biomedicine. Our evaluation of 32 models demonstrates substantial limitations: the strongest model achieves only 75.3% Jaccard Index and 41.8% exact match accuracy. We identify three systematic biases underlying these failures: (i) unselection bias: models systematically avoid certain correct answer choices; (ii) speculation bias: models include incorrect answers when uncertain; and (iii) count bias: models consistently underpredict the number of correct answers. To address these limitations, we propose Choice Funnel, a decoding strategy that combines token debiasing with adaptive thresholding and abstention handling to guide models toward complete and accurate multi-answer selections. Choice funnel improves the accuracy of the exact match by up to 29% while reducing the inference cost by more than 64% compared to the existing approaches. We release SATA-Bench and Choice Funnel to encourage the development of LLMs capable of robust decision-making in realistic multi-answer scenarios.
@@ 1 Introduction
Figure 1: Representative example of an LLM failure on a SATA (Select All That Apply) question. Models often miss valid answers due to unselection, count, and speculation biases. Gemini speculates in this question while GPT-4o underselects. Other models may have unselection bias over C.
Large Language Models (LLMs) have demonstrated remarkable capabilities across diverse natural language processing tasks, with multiple-choice question answering becoming a standard evaluation framework (Pezeshkpour & Hruschka, 2024; Zheng et al., 2024). However, current benchmarks assume a single correct answer per question, even though many applications require multiple valid responses, and because they rely on binary scoring that does not penalize speculation, they inadvertently encourage hallucination (Kalai et al., 2025). Consider content moderation systems that must flag posts for several policy violations simultaneously, medical diagnosis tools that identify co-occurring conditions, or legal research platforms that classify documents under multiple relevant statutes. These scenarios represent Select All That Apply (SATA) tasks, where success depends not on choosing the single best option but on accurately identifying the complete set of correct answers. Despite their prevalence in real-world applications, SATA tasks remain underexplored in LLM evaluation, leaving a gap between benchmark performance and practical utility with direct implications for trustworthiness and safety. Existing evaluations overestimate model reliability by rewarding speculation, whereas SATA-specific metrics directly penalize speculative behavior.
To address this gap, we introduce SATA-Bench, a comprehensive benchmark containing over 10,000 human-validated questions across six domains: reading comprehension, toxicity detection, news categorization, biomedicine, legal classification, and event analysis. Unlike existing multi-label classification datasets that often include dozens of possible labels and assume bag-of-words features, SATA-Bench provides natural-language multiple-choice questions with 3–15 options and 2–10 correct answers, together with metrics that evaluate option-order effects, abstention behavior, and other phenomena unique to LLMs.
Our evaluation of 32 state-of-the-art models (including both proprietary LLMs and open-source alternatives) reveals substantial limitations in multi-answer reasoning. Even the best-performing model achieves only 41.8% exact match accuracy, missing the full correct set in nearly 60% of questions. Figure 1 illustrates a representative failure where models correctly identify some valid answers but systematically avoid others. We identify three systematic biases [^1] We use the term bias to highlight systematic tendencies in prediction (See Appendix P.2 for mathematical definitions), not socioeconomic or demographic bias underlying these failures: unselection bias, where models consistently avoid certain answer positions regardless of content; count bias, where models underestimate the total number of correct answers; and speculation bias, where models include incorrect options when uncertain rather than abstaining (Kalai et al., 2025). To mitigate these issues, we propose Choice Funnel, a decoding algorithm that combines token debiasing, adaptive thresholding, and abstention handling. Beyond evaluation, SATA-Bench serves as both a benchmark and a diagnostic platform, revealing systematic failure modes and enabling algorithmic advances such as Choice Funnel.
@@
    SATA-Bench Data Curation: We curate a high-quality, diverse benchmark dataset explicitly designed to challenge LLMs on multi-answer tasks. SATA-Bench contains more than 10K human-validated questions with multiple domains, varying difficulty levels, multiple correct answers, and carefully constructed distractors. In addition, we provide readability, confusion, and similarity analyses to ensure clarity, diversity, and task complexity across six domains.
@@
    Comprehensive Evaluation: We conduct the largest-to-date evaluation of 32 proprietary and open-source LLMs on SATA questions, revealing that even the strongest models achieve only 41.8% exact match accuracy and 75.3% Jaccard Index.
@@
    Bias Diagnosis: We identify and formalize unselection, count, and speculation biases as obstacles to solving SATA questions, and introduce multiple metrics to evaluate these biases.
  4. 4.
    Choice Funnel Algorithm: We introduce a decoding strategy that jointly mitigates these biases through token debiasing, adaptive thresholding, and abstention handling, improving exact match accuracy by up to 29 percentage points while reducing inference cost by 64%.
@@ 2 SATA-Bench Data Curation
Our objective is to develop a dataset that spans diverse tasks and domains while providing sufficient challenge to reveal differences in LLM capabilities. The curation process consists of three stages: (i) selecting source datasets, (ii) transforming them into SATA format, and (iii) filtering questions for readability, diversity, human validation, and clarity (see Figure 3). We curated SATA-Bench to include tasks in Reading Comprehension (Khashabi et al., 2018), Text Classification (News (Padmanabhan et al., 2016), Events (Event-Classification, )), and Domain Understanding (Toxicity (Gehman et al., 2020), Biomedicine (PubMed-MeSH, 2021), Laws (Chalkidis et al., 2019)). Detailed dataset descriptions are provided in Appendix A.
@@ 2.1 SATA Transformation
We convert each item to a SATA item by first gathering the text, gold labels, and option count. We then enforce an option-to-answer ratio of 2–3 to maintain consistency and difficulty (Thompson & Giffin, 2021). Next, we set $k$ to the number of correct answers $c$, construct the option set with the $c$ gold choices plus kck-c distractors sampled from the pool, and finally shuffle the options to mitigate position and label bias.
Table 1: Compared to prior benchmarks (Kalai et al., 2025), SATA-Bench penalizes speculation, spans multiple domains, uses non-binary metrics, and includes multi-stage human annotations. Penalizing speculation means wrong answers receive lower scores than abstaining. Jaccard Index penalizes speculation: if ground truth is $A,B$ and model predicts $B,C$, $JI(JacardIndex)=0.33$. if it does not speculate and predicts $B$, $JI=0.5$. Thus, this scoring scheme gives a lower score to LLMs that speculate when uncertain.
Benchmark Scoring method Binary Penalizing Human # Domains
grading speculation labeling
GPQA Multiple-choice accuracy Yes None Yes 3
MMLU-Pro Multiple-choice accuracy Yes None Yes 57
IFEval Programmatic instruction verification Yesa None No 1
Omni-MATH Equivalence grading* Yes None Yes 1
WildBench LM-graded rubric* No Partialc Partial Varied
BBH Multiple-choice / Exact Match Yes None Yes 23
MATH Equivalence grading* Yes None Yes 1
MuSR Multiple-choice accuracy Yes None Yes 1
SWE-bench Patch passes unit tests Yes None No 1
HLE Multiple-choice / equivalence grading* Yes None Yes $10+$
SATA-Bench Jaccard Index / Exact Match Partialb Yes Yes 6
* Grading is performed using language models, hence incorrect bluffs may occasionally be scored as correct.
a  IFEval aggregates several binary rubric sub-scores into a composite score.
b  Jaccard Index and Precision are not binary grading.
c  Grading rubric (1-10 scale) may award hallucinated responses.
@@ 2.2 Question Filtering
Initial Filtering. To clean the original source data, we eliminated questions with fewer than ten words (Sanderson, 2010; Karunarathna et al., ). To ensure each question is understandable and solvable, we excluded those containing ambiguous, vague, or subjective terms (Moore et al., 2024). We also removed contaminated questions to reduce memorization risk, following (Li et al., ) (details in Appendix B.1).
Readability. To ensure SATA-Bench questions are both understandable and challenging, we assessed readability using the Flesch Reading Ease (FRE) score (Flesch, 1948) and the Gunning Fog Index (GFI) (Gunning, 1952). We retained questions with an FRE score between 20–100 and a GFI score between 6–17, corresponding to 6th-grade through graduate-level difficulty (Kincaid et al., 1975; Gunning, 1952). This step removed unclear or trivial questions while preserving a broad difficulty range. [^2] We additionally computed four other readability measures—Flesch-Kincaid Grade Level (FGL) (Kincaid et al., 1975), Automated Readability Index (ARI) (Kincaid et al., 1975), and Dale–Chall Readability (DCR) (Dale & Chall, 1948)—which are included in the released dataset.
Question Similarity. To avoid redundancy, we measured cosine similarity between TF-IDF representations (Sparck Jones, 1972) of all question pairs, following (Zhu et al., 2021). We removed questions with at least 80% similarity. We also performed statistical analysis (Appendix B) to confirm the consistency of our label design.
Confusion Score. SATA difficulty is closely tied to the similarity between correct answers and distractors. We quantified this by computing semantic similarity using ST5-XXL (Ni et al., 2021), which performed best in (Muennighoff et al., 2022a). To balance difficulty, we binned questions into 10 groups by confusion score and sampled 50–300 records from each bin, ensuring SATA-Bench covers a wide difficulty spectrum. Figures 5 and 7 show the distribution of confusion scores before and after filtering, as well as breakdowns by source dataset.
Human Validation. Human evaluation proceeded in two stages. First, annotators identified and removed questions containing ambiguous content (Appendix B.2), producing a by-product dataset of 9.5K pre-annotation questions. In the second stage, three annotators reviewed all remaining questions to correct labeling errors. Questions without unanimous agreement were excluded (Appendix B.4). As the outcome, the final release includes a 1.47K evaluation set, and overall statistics of SATA-Bench appear in Table 5.
@@ 2.3 SATA-Bench Characteristics
SATA-Bench has the following characteristics: (i) granular grading: Multiple correct answers provide a finer understanding than binary true/false; (ii) diversity: the dataset spans both knowledge-based and reasoning-driven tasks; (iii) human validation: all items are manually reviewed for clarity and correctness, and readability scores ensure coverage from 6th grade through graduate level, with ambiguous or trivial questions removed; (iv) challenging: 76% of questions fall within the standard FRE range (60–70), the average GFI corresponds to 13th grade (first-year college), and correct answers and distractors have a mean semantic similarity of 0.24 (skewness = 1.8), clustering around 0.22 with a long tail of harder items (Figure 2).
Figure 2: SATA-Bench Evaluation Dataset Overview. SATA-Bench covers a diverse set of topics and achieves a balance between readability and difficulty (measured by confusion score). d1: Reading Comprehension, d2: Toxicity, d3: News, d4: Biomedicine, d5: Laws, and d6: Events.
@@ 3 Experiments
This section presents the experiments conducted to assess the capabilities of LLMs on SATA questions on evaluation set. Our benchmark covers 18 proprietary and 14 open-source models (see Table 8 for details). Because the benchmark spans diverse domains, we adopt a zero-shot evaluation protocol. The system prompt specifies that each question has at least two correct answers, and we instruct the LLM to output the results in JSONL format (Intelligence, 2024; Zhou et al., 2023). Furthermore, we employ a CoT prompting strategy following (OpenAI & el at, 2024). We then extract answers from the JSONL output using exact and fuzzy match. For cases where JSONL extraction fails (fewer than 3%), we use Claude 3 Haiku and human labelers to recover the correct options. However, for smaller models, the JSONL extraction fails in more than 5% of the cases, making this method less reliable. In these cases, following (Hendrycks et al., 2021), we omit CoT and instead rank options using the probability of the first output token. To calibrate thresholds, we hold out 100 randomly sampled instances from the benchmark and tune each model for the optimal Jaccard Index (Bogatinovski et al., 2022). We then select all options with probabilities above this threshold. This probability-based method applies only to models with accessible token likelihoods. Finally, we also evaluate the performance of non-expert humans on the benchmark (Appendix E).
Table 2: Performance comparison of 32 different LLMs across various metrics on SATA-Bench. We highlight the best (bold) and second-best (underline) values. Columns labeled [($\uparrow$)] indicate higher-is-better; columns labeled [($\downarrow$)] indicate lower-is-better. Models with explicit reasoning capabilities are highlighted in italic. All numeric values are rounded to two decimal places. We retrieve exact labels for models evaluated using Inference-Based Retrieval + CoT prompting. For models evaluated under Probability-Based Retrieval, we select labels based on token probability thresholds.
@@
O3 73.91 31.58 41.77 87.50 0.38 6.79 0.06 -0.39 0.94 46.12
GPT 4.1 75.23 40.37 40.49 85.52 0.13 5.98 0.06 -0.04 0.85 45.52
GPT-OSS 120B 74.28 37.53 40.29 86.28 0.19 6.31 0.07 -0.16 0.84 47.57
Grok 3 Think 74.40 43.10 39.71 83.93 0.30 6.26 0.07 0.06 0.93 44.24
GPT 4 74.11 38.42 39.47 85.90 0.21 6.63 0.06 -0.20 0.82 46.61
Claude 3.7 Think 70.96 35.16 37.92 85.03 0.46 18.77 0.34 -0.32 0.87 44.48
Claude 3.7 70.98 33.10 37.82 85.35 0.49 6.59 0.25 -0.43 0.93 43.58
Claude 3 Sonnet 70.72 38.81 36.49 84.58 0.36 7.37 0.07 -0.35 0.83 48.00
Geimini 2.5 Think 72.58 42.16 36.46 84.58 0.12 4.76 0.06 -0.01 0.88 43.76
Claude 3.5 Haiku 71.12 50.01 35.89 80.26 0.33 7.31 0.35 0.18 1.01 42.61
Claude 3 Haiku 70.63 40.84 35.64 83.59 0.42 6.24 0.07 -0.22 0.85 47.15
Claude 3 Opus 70.15 34.17 35.59 86.97 0.62 8.26 0.07 -0.52 0.93 44.36
Gemini 2 Flash 70.71 40.79 34.60 85.01 0.17 6.14 0.06 -0.23 0.91 39.94
GPT 4.1 mini 69.90 37.31 33.46 86.05 0.30 6.69 0.06 -0.39 0.97 38.61
Nova Pro 68.92 31.64 32.95 87.37 0.52 7.92 0.07 -0.55 1.01 39.27
Claude 3.5 Sonnet 67.15 34.25 32.22 87.57 0.43 8.41 0.09 -0.46 1.06 38.55
Llama 3.1 405B 67.18 35.06 30.17 86.24 0.33 6.90 0.45 -0.39 1.02 36.30
Nova Lite 63.75 39.88 29.11 82.51 0.52 9.12 0.45 -0.51 1.17 37.39
Deepseek R1 64.49 34.89 28.17 84.62 0.94 17.44 0.03 -0.57 1.13 33.52
GPT-OSS 20B 60.73 40.90 27.35 80.90 0.77 11.05 0.10 -0.53 1.45 31.80
Mistral Large V2 57.16 27.23 22.83 88.20 1.33 10.89 0.12 -1.10 1.47 27.27
Qwen Plus 55.74 24.03 21.12 88.54 2.24 10.72 0.11 -1.18 1.43 24.85
Nova Micro 55.77 29.28 18.37 86.06 1.84 11.10 0.27 -1.09 1.41 24.30
Llama 3.2 90B 55.78 23.81 18.30 89.56 1.84 11.10 0.27 -1.09 1.41 24.30
Llama 3.1 70B 55.59 23.92 17.94 89.56 1.81 10.06 0.10 -1.12 1.48 22.12
Non-expert Human 45.02 17.93 60.62 1.46 15.32 1.46 -0.6 1.44 34.12
@@
Mistral 8B 46.63 32.21 14.73 81.46 11.42 19.47 1.27 -1.35 1.95 21.01
Llama3 8B 43.64 30.06 13.82 80.30 12.09 17.85 1.09 -1.59 1.88 22.00
Bloomz 7B 41.15 57.76 11.27 66.09 20.62 29.00 1.51 -0.87 1.71 20.09
DeepSeek R1 Distill 8B 40.02 45.33 8.85 72.20 13.38 21.62 1.14 -1.29 1.75 20.42
Qwen2.5 14B 37.58 17.27 6.30 87.84 21.01 18.02 1.06 -2.24 2.26 11.93
Phi3 7B 34.57 17.64 2.97 87.25 23.22 18.57 1.22 -2.33 2.35 7.22
Phi4-mini-reasoning 29.69 26.73 2.12 77.98 21.62 13.90 1.59 -2.37 2.39 7.35
3.1 Evaluation Metrics
Evaluation of SATA question responses requires metrics that capture partial correctness , penalize inappropriate selections, and identify bias. We organize our evaluation into three categories: performance metrics that measure correctness and speculation, selection bias metrics that quantify positional preferences, and count bias metrics that assess quantity prediction accuracy. Detailed explanations for all metrics appear in Appendix F.
Performance and Speculation Bias Metrics. We employ four metrics to assess answer correctness (Tarekegn et al., 2024a). Jaccard Index (JI) measures the intersection-over-union between predicted and gold labels, providing credit for partial matches. A low JI also reflects limited overlap between predicted and gold labels, indicating speculation bias. False Positive Rate (FPR) measures the proportion of questions where models select any incorrect option, directly quantifying speculation bias. Exact Match (EM) requires the predicted set to exactly match the gold set, representing the most stringent evaluation criterion. Mean Average Precision (Precision) evaluates the fraction of selected answers that are correct, which also gives credits to partial correctness.
(Un)selection Bias Metrics. To characterize positional preferences, we measure models’ tendencies to favor or avoid specific option positions. We use RStd (Zheng et al., 2024) and RSD (Croce et al., 2020; Reif & Schwartz, 2024) to quantify selection bias toward particular option IDs. Additionally, we introduce Selection Probability Divergence (SPD) to measure unselection bias—the systematic tendency to avoid certain options regardless of content (detailed in Appendix G).
Count Bias Metrics. Models often select fewer options than warranted, necessitating specialized metrics for quantity assessment. We measure: (i) mean signed difference between selected and correct counts (CtDif), where negative values indicate under-selection; (ii) mean absolute difference (CtDifAbs) to quantify magnitude regardless of direction; and (iii) percentage of cases with exact count matches (CtAcc) to assess quantity prediction accuracy.